OpenStudy (anonymous):

Question 5 (Multiple Choice Worth 2 points) Simplify square root of -72 –10 –10 10 10

OpenStudy (anonymous):

–10 –10j 10j 10 is the answers

OpenStudy (anonymous):

opps its the square root of -100

OpenStudy (anonymous):

\[\sqrt{-72}= \ \ \sqrt{72 i ^2}=\ \ \sqrt{3^{2} *2^{2} * 2 i ^2}= \ \ 6i \sqrt{2 }\]

OpenStudy (anonymous):

Cant be any of those because 72 is negative

OpenStudy (anonymous):

10i